4. How much sucrose should be weighed out to prepare 80 ml of a 1. 0 M solution?

  1. A student weighs out the correct amount of sucrose, places it in a beaker, measures 80 ml of dH2O with a graduated cylinder and adds the dH2O to the beaker. The solution is mixed until all the sucrose is dissolved. What did the student do wrong?
  1. If the final volume of the solution prepared was 95 ml, what is the actual molarity of the solution?
Answer #1

The molar mass of sucrose is 342.3 g/mol

1.0M solution is 1 mole of a substance dissolved in 1000mL solvent, in case of sucrose when 342.3 g is dissolved in 1000mL it will give 1.0M solution.

However, we need an 80mL solution.

So the mass of sucrose needed = (80*342.3)/1000 = 27.38g

Thus 27.38 g of sucrose is needed to be dissolved in 80mL to prepare 1.0M solution.

a) Graduates cylinder may have an error in the calibration and the marks are not well standardized. Hence student made a mistake by taking 80mL water in the cylinder. Instead, the student should have added 27.38 g of sucrose in a standard volumetric flask and then add the water to dissolve it completely in the minimum amount of water to give a clear solution. Further, water should have been added to the mark.

b) Now we have 27.38g in 95mL (i.e. 95*10^-3 L).

The number of moles of sucrose = 27.38 g / 342.3 g/mol = 0.0799 mol

Molarity = moles / L = 0.0799/(95*10^-3) = 0.84 M

Thus the actual molarity of a solution is 0.84 M.